OpenStudy (anonymous):

solve for m 5/11=m/418

OpenStudy (anonymous):

I know the answer, but what do you think the answer is?

OpenStudy (anonymous):

no /there are no answer chioces

OpenStudy (anonymous):

Okay. Well m=190.... Here is how I go the answer.

OpenStudy (jenma):

190M

OpenStudy (anonymous):

Jenma really

OpenStudy (anonymous):

im helping her.

OpenStudy (paki):

@Ddaddysgirl14 try to solve by leaving "m" alone.... @vbroberson please avoid posting direct answers....

OpenStudy (anonymous):

5/11=m/418 Step 1: Cross-multiply. 5/11=m/418 (5)*(418)=m*(11) 2090=11m Step 2: Flip the equation. 11m=2090 Step 3: Divide both sides by 11. 11m/11=2090/11 m=190 Answer: m=190

OpenStudy (jenma):

5/11 = m/418 then you divide 5 to 418 and 11 to m then you got 2090/11m and divide it, you will get the answer ;) i hope this help

OpenStudy (anonymous):

@paki back off... I wasn't posting direct answers... for your information, i was about to exp;ain how I got the answer sobefor eyou judge let one other type...

OpenStudy (anonymous):

@Jenma Chill!!! I just told her.

OpenStudy (anonymous):

ok thx so much can u help with a couple more? @vbroberson

OpenStudy (anonymous):

I'll try

OpenStudy (paki):

@vbroberson try to be nice... please

OpenStudy (jenma):

I'm trying to help her too.

OpenStudy (anonymous):

@paki Okay.

OpenStudy (paki):

thanks for positive feedback.... have a nice day...

OpenStudy (paki):

@Jenma nice work.... like it...

OpenStudy (jenma):

thank u @paki i hope @Ddaddysgirl14 will understand my answer :3

OpenStudy (anonymous):

@Daddysgirl14 Do you have another question?

OpenStudy (misssunshinexxoxo):

@vbroberson very amazing job please no direct answers but great work:) cheers all

OpenStudy (anonymous):

How much interest will be paid in 2 years for a of $1500 at 18.25% simple interes?

OpenStudy (anonymous):

Why does everyone say i'm saying direct answers.. that wasn't me that was jenma, I said the aswer, but then i said I was going to solve it.. jenma was the one who said the answer, and then start trying to explain it when I was helping her.... gosh let me get off here cuz im getting a bit mad. @Ddaddysgirl14 I might not be able to help you on this one it's a bit hard.

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@paki can u help me on a few more and this one/ ^

OpenStudy (jenma):

oh so you're guilty with that. well i'm just trying to help her too and not only one person who can answer it but everyone will.

OpenStudy (anonymous):

stop fighting yall did great helping now see bout this one ^

OpenStudy (anonymous):

@Jenma SHUT THE HECK UP BECAUSE YOUR TICKING ME OFF your still on here arguing about something so stupid, just shut it.

OpenStudy (anonymous):

so what if he gave ' direct answer' at least he still helped

OpenStudy (anonymous):

ok can someone at least hlelp me with the one after the solve for m one^

OpenStudy (anonymous):

hey vbobroson look under your profile and what i wrote bout u testimonial

OpenStudy (anonymous):

how much interest will be paid in 2 years for a loan of 1500 of 18.25% simple interest @peachpi and others

OpenStudy (teddyiswatshecallsme):

@vbroberson Chill bruh. You're the one freaking out dude. XD

OpenStudy (anonymous):

@Teddyiswatshecallsme I'm chill now.

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

ABC and DEF are similar triangles. what is the length of EF A.20cm B.36cm C.42cm D.45cm

OpenStudy (anonymous):

help?!?